General Meeting Summary 2/8/17
Jump to navigation
Jump to search
AEROGEL DETECTOR PERFORMANCE (Arthur)
- Aerogel detector performance with cosmics (JLab passwored required): https://logbooks.jlab.org/entry/3456299
- one channel not counting - problem is likely a cable or connector
- signal attenuation when going to counting house - on floor have 1 V signals
- Hall C detector Wiki: https://hallcweb.jlab.org/wiki/index.php/Wednesday%2C_February_07%2C_2017
- CUA main priority is to check if see signals (SEP), identified potential cabling issues and working on fixing them
- software: can adapt code from standalone version to general Analyzer
- Aside: before final installation need to switch tray to SP-30
TEMPERATURE DEPENDENCE OF PBWO4 LIGHT YIELD (Salina)
- Performed measurements at 18, 15, 10, and 5 degC so far - observed that light yield increases with decreasing temperatures consistent with literature
- trying to reach sub-zero temperatures
- need to check performance/attentuation of optical grease (Bicron?) at these temperatures - what did PANDA use?
- simple test: measure PMT response with LED and layer of optical grease at different temperatures
- practical consideration: warm up crystal before removing it from PMT
- Temperature gradient when cooling: currently waiting ~24 hours for crystal to equilibrate before taking measurements
- notice that on 2/7 freezer temperature went up to 15 degC for HIPS2017 tour
STATUS OF NPS PROTOTYPES (Arthur)
- 1x1 single crystal has been tested
- 3x3 needs some re-assembly - 8 crystals available for it
- Re-cap of Timelike Compton Scattering discussion from HIPS2017
- Perform simulations to make projections relative to other experiments on market - may ignore polarization
PIXEL DETECTORS (Rachel)
- Submitted Quarknet proposal for high school teachers to spend time at Fermilab and CERN learning about pixel detector research
- Detector installation is going well - discussion about additional developments
NEXT MEETING: WED 15 FEBRUARY AT 10:00 AM (EST)